OpenStudy (anonymous):

1+2+5+10 is supose 2 equal 17 some how

OpenStudy (anonymous):

what does the problem say?

OpenStudy (zepp):

1+2 = 3 3+5 = 8 8 + 10 = 18, not 17

OpenStudy (anonymous):

exactly

OpenStudy (lgbasallote):

but if you make it 1 x 2 + 5 + 10 then it's 17 :P

OpenStudy (zepp):

No lgbariddles pls :(

OpenStudy (anonymous):

1 2 5 10 your suppose to move it to the other side and equal 17

OpenStudy (lgbasallote):

1 x 2 + 5 + 10 = 17 mwahahaha

OpenStudy (lgbasallote):

just rotate the addition symbol by 90 degs

OpenStudy (zepp):

45 degrees *

OpenStudy (anonymous):

i need help please

OpenStudy (accessdenied):

90 degrees is still a + in real arithmetic, though: 

OpenStudy (lgbasallote):

whatever lol i hated trig mwahahah

OpenStudy (lgbasallote):

oh wait..that's algebra o.O

OpenStudy (lgbasallote):

scratch that..i hated graphing :p

OpenStudy (zepp):